7.3 kJ of energy required to melt 0.65 mol of ethylene glycol at its melting point. Calculate the 

molar enthalpy of fusion of ethylene glycol and the energy absorbed.


1
Expert's answer
2021-05-09T00:00:23-0400

Standard enthalpy change of fusion of ethylene glycol "(Δ_{fus}H_o)= 9.9\ kJ/mol"


To melt 0.65 mol,

Q = ∆H × n = 9.9 × 0.65 = 6.435 kJ

"\therefore 6.435\ kJ" is absorbed to melt the 0.65 mol of ethylene glycol.